# Evaluating Compiler Optimization Impacts on zkVM Performance

### v1

@@ 6.2. Lessons Learned and Discussion
Optimized code may uncover bugs in zkVMs. In RQ2 (Section 4.2), autotuning a specific program exposed SP1 bugs, including a security-critical case where the zkVM silently aborted mid-execution yet still produced a proof that passed verification. Such issues show that optimized code can stress parts of a zkVM that unoptimized inputs miss. These bugs highlight the need to improve zkVM reliability, which is also exemplified by bug bounty programs for many zkVMs (HackenProof, Inc., 2025; code4arena, 2025). Inspired by compiler testing (Yang et al., 2011; Le et al., 2014), zkVM testing could use optimized vs. unoptimized runs as a test oracle, that is, two equivalent versions should produce identical results, with any divergence indicating a bug.
@@ 8. Conclusion
We presented the first systematic study of how compiler optimizations affect two production RISC-V zkVMs. Most passes help, yet gains are modest and a few are harmful (e.g., licm). We established a strong correlation between cycle count and proving time, enabling cycle-count-guided autotuning that yields up to $2.2\times$ proving speedups beyond -O3, and we prototyped zkVM-aware LLVM changes that further improve the performance of zkVMs.
Future work can extend this study to additional zkVMs and proof systems, unifying cost metrics (e.g., prover-gas-like models), and exploring superoptimizations for zkVMs. These compiler-centric directions can push zkVM proving performance beyond what proof-system research and engineering alone can deliver.
@@ Appendix A Differences between zkVMs and Native Execution
Instruction latency. On traditional architectures, instruction latency can vary significantly depending on the instruction type, cache hits/misses, and other micro-architectural effects (Fog, 2022). In zkVMs, however, most instructions have a uniform cost (ris, 2025a), with most instructions taking exactly the same number of cycles.
Memory access. Memory access is relatively cheaper on zkVMs. For example, on RISC Zero, memory access costs only one cycle if the page is already paged in. Page-ins and page-outs on RISC Zero are, however, more expensive, taking around 1130 cycles on average (ris, 2025a). zkVMs also do not have any caches making memory access always synchronous.

### v2

@@ 6.2. Lessons Learned and Discussion
Combined cost metrics. Our analysis in Section 5.1 establishes that zkVM performance is primarily driven by dynamic instruction count. Paging cycles provide complementary information. This motivates the use of combined cost models that integrate multiple components into a single performance predictor. While this two-factor model explains the majority of performance variation (Table 2), some effects remain unexplained, such as SP1’s proof sharding overhead in Section 6.1. Future work could refine the model by incorporating more zkVM-specific factors (e.g., use of precompiles), or learning workload-specific weights (e.g., memory-heavy vs. computation-heavy programs).
@@ 8. Conclusion
We presented the first systematic study of how compiler optimizations affect two production RISC-V zkVMs. Most passes help, yet gains are modest and a few are harmful (e.g., licm). Autotuning yields up to $2.2\times$ proving speedups beyond -O3. We established a strong correlation between dynamic instruction count and zkVM performance and, based on this insight, prototype zkVM-aware LLVM changes that further improve zkVM performance.
Future work can extend this study to additional zkVMs and proof systems, unifying cost metrics, and exploring superoptimizations for zkVMs. These compiler-centric directions can push zkVM proving performance beyond what proof-system research and engineering alone can deliver.
@@ Appendix A Differences between zkVMs and Native Execution
Instruction latency. On traditional architectures, instruction latency can vary significantly depending on the instruction type, cache hits/misses, and other micro-architectural effects (Fog, 2022). In zkVMs, however, most instructions have a uniform cost (ris, 2025), with most instructions taking exactly the same number of cycles.
Memory access. Memory access is relatively cheaper on zkVMs. For example, on RISC Zero, memory access costs only one cycle if the page is already paged in. Page-ins and page-outs on RISC Zero are, however, more expensive, taking around 1130 cycles on average (ris, 2025). zkVMs also do not have any caches making memory access always synchronous.